OpenStudy (anonymous):

John has read the first 114 pages of a novel. He has read 3 pages less than one-third of the novel. Write an equation to describe the total number of pages p in the novel, and find the value of p. Hellpppppp

OpenStudy (bohotness):

less than means you are going to subtract

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{\dfrac{p}{3}-3=114}\end{align}\)

OpenStudy (bohotness):

114 is equal to the number of pages he has read so far let x represent the number of total pages in the book set 114 equal to what 114 is as a part of the book(3 pages less than 1/3 the number of pages in the book)

OpenStudy (anonymous):

114+3+2/3P=P 114+3+117 Since 117 is 1/3 of the book, multiply it by 3 117*3=351